Start-up Observing Night 1 - 20170902 UT

Observers: JHill, OKuhn, BRothberg, DMiller (Tucson) SWalsh (LBTO)
Telescope Operator: SAllanson (LBTO)
AO Support: GTaylor
Telescope Support: KNewton (LBTO)
Software Support: CBiddick, KSummers (Tucson)
Instrument Support: ESolheid (LBTO)
Engineering Support: TMcMahon

Summary

LBC had some problems, so we switched to doing AGw1/2 wake-up at the start of the night.

TCS Build 2017A has been confirmed on sky for pointing/guiding/active optics/adaptive optics. Ran LBC hijacked NS presets toward the end of the night. Works fine in regular NS mode, but some issues when updating from ephemeris (IT 6743, IT6746, and IT 6750 - tomorrow).

AGw1 and AGw2 have been woken up. The probes have been aligned to the rotator center although SX couldn't quite get there due to limited probe travel (more or less a feature).

SX and DX M3 positions for LUCI have been adjusted to minimize the wobble of the S-H pupils.

LBC checkout tasks that did not require guiding done: rotator center, dohybrid, dofpia, superfoc, set of exposures through all filters, dithering. But because trackers did not come up (IT 6741) we could not run a science-like OB. Noticed speckled arcs near bright stars on R and r images taken with LBCR, from a laser (IT 6744 - later found to be the 4D laser). Tested NS observing with varying rates. OK, but some image jumps (related to this or not? IT 6746).

Preparations

00:21 Olga and Elliott running testrotator on LBC-Red for cable chain checkout in parallel for LBC fills. Problems getting LBC to start-up properly. Called Kellee. (You must run testrotator as root, and misc other LBC crankiness.)

00:40 Rotator and filter wheel check complete on both LBCs. Problems getting RedSci and RedTech to come up - not talking to RedSci with either fiber parity. Change plans to work on FBG AGw alignment.

NSIGUI won't retrieve ephemeris file on obs1. couldn't execute "ftp": no such file or directory It works OK on obs2.

02:05 Last RedSci problem was that power to controller was off due to using engineering interface.

Details

Open

01:58 First attempt to open shutter doors. One door not reading correctly on ECSGUI. Both FBG rotators showing LOCKOUT, but wasn't DaveT just using them?

SOME WIKI EDITTING DISASTER WIPED OUT A BUNCH OF MY NOTES HERE. WE FOUND STARS AND COLLIMATED AT LFBG/RFBG.

Alignment of AGW1 / AGW2 guide probes to rotator center (BGTechProcedures)

John doing DX, Olga and Barry doing SX.

DX guide 00002 centered at 126.5,114 (binned)
253,228 (x2)
270,243 (detsec)
DX Hotspot is 275,243

03:20 Adjust DX X by 0.15mm * 68 cts/mm = +10 counts in homeoffset.

03:20 Adjust SX X by 0.345 mm * 68 cts/mm = +23 counts in homeoffset
Adjust SX Y by 0.216 mm = +216 microns, but had to reduce that to +100 microns due to probe travel
Note: We (JMH) UNDID this +100 micron correction on 20170904 UT, as it was still causing problems with the guide probe being sticky as it got near the front limit. On 20170905 UT we went back to the full +216 microns as the stickiness seems not to depend on the exact position.

3:50 Checking the corrections

DX guide 00003 centered at 123.5,116 (binned) but seeing is substantially worse (1.8 arcsec)
247,232 (x2) 264,247 (detsec)
Aarghhh, John went the wrong direction. Now fixed to -10 counts.

FBG Pupil wobble alignment of M3

DougM has joined us for this test.

04:06 Doug struggling for a few minutes with unexpected comma at the end of wfs filename string.

04:46 LFBG align T/T of Tertiary - pupil_collect - pupil_process_set, pupil_04:26:09.dat

  IE CA Pupil dX pixels Pupil dY pixels

| M3 Tip | +200 | +7.5 | +15.0 | +1.38 | +1.86
M3 Tilt -300 +16.0 -8.5  
SX M3
04:48 Tip=+200 Tilt=0  dIE=+7.5 dCA=+15
04:53 Tip=-200 Tilt=-300   dIE=+16 dCA=-8.5 (for the tilt)
05:07 Tip=-200 Tilt=-600
05:15 Tip=-200 Tilt=+600
05:22 Tip=-400 Tilt=+300
05:28 Tip=-600 Tilt=+300
06:25 Tip=-600 Tilt=0   (Absolute Tip=-580 Tilt=-160)  ADOPTED
DX M3
06:10 Tip=+200 Tilt=0  dIE +7.5 dCA=-15   (CA corrections have opposite sign)
06:18 Tip=0 Tilt=+600  dIE -32 dCA -17
06:28 Tip=-200 Tilt=+300 (Absolute Tip=-1200 Tilt=+400) ADOPTED

06:30 SX and DX M3 alignment completed (new values not installed in the collimation files until the afternoon of 03-Sep-2017)

SX Focal Station Angle =+26
For SX Tip dIE=Tip/12*sin(FSA) dCA=Tip/12*cos(FSA)
For SX Tilt dIE=-Tilt/18*cos(FSA) dCA=-Tilt/18*sin(FSA)

DX Focal Station Angle=-26
For DX Tip dIE=-Tip/12*sin(FSA) dCA=-Tip/12*cos(FSA)
For DX Tilt dIE=-Tilt/18*cos(FSA) dCA=Tilt/18*sin(FSA)

SX AO Checkout

06:50 Doug attempting to close SX (and DX) AO loops.

07:20 SX AO loop closed, so switch over to LBCs.

LBC Wake-up and Collimation

We have LBCs run up but with both red and blue trackers unchecked. These are not working, see IT #6741.

07:48: Slew to Galadi Field copoint (will use field for rotcen check).

07:50: dohybrid: the wrs part worked fine, but fpia threw an error on b075734 (due to clouds?):
Reading "lbciaCoeffs_red.dat" file ... Sending data to TCS ... Success

Checking for Red Convergence
Limits:    250.0   250.0   250.0   400.0   400.0   250.0    50.0
2017-09-02
Processing: /newdata/lbcb.20170902.075734.fits
Searching for pupils... 
Computing aberrations using 5 pupils
Focus Offset = -0.80000001  Filter = V-BESSEL
BlueZ22Zero:   1.600
RedZ22Zero:   1.650
% POLY_FIT: Singular matrix detected.
% Execution halted at: TRYCENTROID        22
   /home/lbceng/WRS/src/trycentroid.pro
%                      FITPUPIL          470 /home/lbceng/WRS/src/fitpupil.pro
%                      CALCABCOEFFS      136
   /home/lbceng/WRS/src/calcabcoeffs.pro
%                      LBCFPIA           284 /home/lbceng/WRS/src/lbcfpia.pro
%                      DOHYBRID          826 /home/lbceng/WRS/src/dohybrid.pro
%                      $MAIN$          
% Program caused arithmetic error: Floating underflow
% Program caused arithmetic error: Floating illegal operand

Try fpia again. Worked fine, converged in 1 iteration.

08:03: Send copoint OB. b 080353 & r 080350
MOUNT: delta_IE = 9.81 arcsec, delta_CA = 0.79 arcsec
M1 SX: dX = 0.28 mm, dY = -0.03 mm, dRX = -0.72 arcsec, dRY = -5.97 arcsec 
M1 DX: dX = -0.35 mm, dY = 0.00 mm, dRX = 0.09 arcsec, dRY = 7.49 arcsec 

08:08: Check rotator center.
galadifield_rot0 pa=0
galadifield_rot180 pa=180 and dec=-90 (last minute edit to dec. Somehow 10 0,0 dithers crept in, stopped OB after a few.)
galadifield_rot90 pa=90 (also had 10 0,0 dithers)
galadifield_rot270 pa=270 and dec=-90
  • Barry determines LBCR rotcen @ 1081, 2913 and LBCB rotcen @ 1029, 2915

08:22: Take superfoc sequences:
  • Same field. Run FPIA (08:23)
  • clouds coming in and out of field. Nearly converged with seeing ~ 1" but last measurement had clouds & larger seeing estimate and may have thrown off result. Also, field is getting very high (dec 30). Going to new field.
  • SA92 field. Run FPIA with 16-sec and then 32-sec to get more pupils (4-5 pupils now, ~same as before).
  • Converged
  • 08:42 started RB_rV_sf7 PA=0
    • Barry measures best-focus for r 0.008 and V -0.0031
  • 08:49 started RB_rV_sf7_pa180 PA=180
    • best-focus for r -0.047 and V -0.020
  • dofpia, /x2 again - blue took a while
  • 09:05 started RB_rV_sf7_pa180
    • best-focus for r 0.0054 V -0.0260
  • followed immediately by RB_rV_sf7
    • best-focus for r 0.00040 V -0.0205

09:23: Take 60-sec images through all filters (like the closed dome test). We saw ghost image or reflection from bright star on red images taken previously. Check whether this is just for r-SLOAN or other filters, too. Also, 60-sec but no tech chips functioning seems to work even when guiding is left ticked. Didn't get very deep F972, Y or V images, but got 60-sec images at I, R, z, r, i and we don't see this feature except at R and r. Laser? (IT 6744)

NS test observations of OSIRIS

09:43: Slew to Al's NS target, we thought. No, now going to his target.

09:47: dofpia. Skies look clear now, although there had been pretty extensive cirrus earlier.

09:54 started non-sidereal science exposures of OSIRIS. This tests TCS handling of quickly changing rates.

10:06: John tried to use new PCS feature that updates ns rates from ephemeris, but it apparently crashed PCS. Steve restarted PCS and we started a new 5-pt dither on the target. See IT 6743.

10:18:40 John tried same thing as before, but it did not crash PCS. Ran new 5-pt dither.

10:29: dofpia, followed by 5-pt dither this time of 300-sec exposures.

11:10: run BI ob. First pair of b r images r 11155, b 11159 have trailing. 60-sec exptime. Jumps. 2nd pair also shows jumps, but 3rd is fine. b 111922 r 111919

11:22: stopped OB that was jumping. Run fpia. Converged in first iteration.

11:26: run OB again. Note that there were quite a few pairs of images, starting with the BI OB so around 11:10UT, which showed image jumps - was this because we were tracking non-sidereally, not guiding, or is there another reason? IT 6746

12:15: saturating on 60-sec exposures. Run sky flats:

* V V at PA=0 and PA=180.

* Tried S z but saturating.

Close

12:38: Closing up. Skies look clear on allsky with some cirrus on the horizon. As far as LBC checking, we did most of the checkout, but the critical test of running a science-like OB with guiding and dithering was not done because the trackers were not working.

12:47 Steve troubleshooting problem with shutter door (doors are closed and dome dark). Taking 25 biases, last red one showed horizontal banding, although I don't think anything was done in the dome. Took 2 more for good measure (look OK) and I will take more biases at the next opportunity.

-- %USERSIG{JohnHill - 2017-09-02}%
Topic revision: r13 - 26 Oct 2017, JohnHill
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback